# Archilles’ Heel in Semi-open LLMs: Hiding Bottom against Recovery Attacks

### v1

@@ C.3 Customization Performance of Models
Table 21: Detailed results of Llama2-70B closed by SCARA on six downstream tasks.
@@
Table 22: Detailed results of Llama2-7B closed by SCARA on six downstream tasks.
@@
Table 23: Detailed results of Mistral-7B closed by SCARA on six downstream tasks.
@@
Table 24: Detailed results of Phi-2 closed by SCARA on six downstream tasks.
@@
Table 25: Detailed results of Phi-1.5 closed by SCARA on six downstream tasks.
@@
C.4 Comparison in closing baselines on llama2-70B
We compare the recovery resilience of SCARA with SAP-DP and Fully-closed as baselines under FT-closed and SEM attack strategies. The evaluation results on sixteen benchmarks are shown in Table 26.
Table 26: Evaluation results of Llama2-70B under FT-closed and SEM attack
FT-closed SEM
SCARA SAP-DP SCARA SAP-DP
@@ C.4 Comparison in closing baselines on llama2-70B
Average Recovery Ratio($\downarrow$) 22.60 21.80 22.40 22.30
C.5 Comparison in Recovery Attack Strategies
In this section, we present detailed evaluation results of the model recovery performance of SCARA under FT-closed and SEM attack strategies across six functionalities used in our experiments. The detailed results under the FT-closed recovery strategy are presented in Table 27. The results under SEM attack strategies are shown in Table 28.
Table 27: Recovery Performance of SCARA under FT-Closed attacks.
@@ C.5 Comparison in Recovery Attack Strategies
Average Recovery Ratio($\downarrow$) 24.80 22.50 23.56 26.97
Table 28: Recovery Performance of SCARA under SEM attacks.
@@
Average Recovery Ratio($\downarrow$) 25.00 22.00 22.10 24.70
C.6 Comparison in Recovery datasets scales
To investigate the impact of attack dataset scales on the efficiency of SCARA, we conduct model recovery attack on the Llama2-7B model using four different attack datasets of varying sizes: 100k, 200k, 300k, and 500k. The evaluation performance under different attack set scales are in Table 29
Table 29: Evaluation Results of SCARA on Llama2-7B under Various Attack Set Scales.
@@ C.6 Comparison in Recovery datasets scales
Average Recovery Ratio($\downarrow$) 25.07 25.03 24.89 25.26 25.48
@@
Resilience Performance. We close same-sized layer sets with different start points, and attack them using FT-all. Specifically, the sets consist of one layer for Llama2-7B (Table 30, Table 31), and two layers for Phi-2 (Table 34, Table 35). We further computed the $\Delta\text{ARR}$ for each closed-source set within Mistral-7B-v0.1 and Phi-1.5 in Appendix B.8. The results for the Mistral-7B-v0.1 model are presented in Table 32 and Table 33. Additionally, the performance outcomes for the Phi-1.5 model can be found in Table 36.
In all the above tables, “Pretrain” represents the model’s original performance without any layers closed-sourced. These columns indicate the index of layers in the model that have been closed-sourced. “*” indicates fully-closed. All evaluation scores are averages from three different seed tests, corresponding to the values 20, 42, and 1234, following the details of the Sixteen Functionality Benchmarks in Appendix B.5.
Table 30: Evaluation Results of Llama2-7B under Different Closed Layers (Part1)
@@
Average Recovery Ratio($\downarrow$) - 30.76 30.67 29.55 30.39 29.99 30.13 31.04 31.90
Recovery Difficulty($\uparrow$) - 11.11 11.27 10.87 10.31 10.83 10.33 10.90 11.11
Table 31: Evaluation Results of Llama2-7B under Different Closed-sourced Layers (Part2). “*” indicates the fully closed-sourced model.
@@
Average Recovery Ratio($\downarrow$) 36.89 43.94 50.98 63.48 68.87 75.35 80.27 86.24 31.39
Recovery Difficulty($\uparrow$) 10.42 9.49 8.86 7.12 6.14 4.72 3.40 3.06 11.19
Table 32: Evaluation Results of Mistral-7B under Different Closed-sourced Layers (Part1)
@@
Average Recovery Ratio($\downarrow$) - 26.38 27.85 27.32 26.25 26.97 26.15 26.24 28.20
Average Recovery Ratio($\downarrow$) - 11.50 11.31 11.48 10.71 10.77 11.44 11.02 10.71
Table 33: Evaluation Results of Mistral-7B under Different Closed-sourced Layers (Part2)
@@
Average Recovery Ratio($\downarrow$) 28.83 28.71 29.84 32.20 42.09 49.64 48.64 64.08 26.05
Average Recovery Ratio($\downarrow$) 11.34 11.11 10.45 10.59 10.23 10.34 9.59 8.53 11.20
Table 34: Evaluation Results of Phi-2 under Different Closed-sourced Layers (Part 1)
@@
Average Recovery Ratio($\downarrow$) - 29.15 55.90 58.99 63.21 71.00 71.97 72.28 76.09
Recovery Difficulty($\uparrow$) - 10.07 7.07 4.95 4.09 3.63 3.31 3.31 3.11
Table 35: Evaluation Results of Phi-2 under Different Closed-sourced Layers (Part2). “*” indicates the fully closed-sourced model.
@@
Average Recovery Ratio($\downarrow$) 72.51 71.58 71.38 74.04 77.22 82.04 85.63 87.87 25.59
Recovery Difficulty($\uparrow$) 3.07 3.29 3.03 3.01 2.70 2.32 1.98 2.13 11.32
Table 36: Evaluation Results of Phi-1.5 under Different Closed-sourced Layers
@@
Average Recovery Ratio($\downarrow$) - 35.02 55.90 61.66 69.14 72.29 72.34 76.80 75.90 81.68 84.25 86.69 89.56 33.75
Average Recovery Ratio($\downarrow$) - 10.08 7.18 4.70 3.50 2.93 2.83 2.53 2.36 2.27 2.16 2.06 2.46 9.33
@@
Table 37: Customization Performance under Different Closed Sets
@@
Closed Layers GSM8K($\uparrow$) Closed Layers GSM8K($\uparrow$)
@@
Fully-Closed 20.32 Fully-Closed 57.77
C.8 Evaluation Results under Different Closed size
In this section, we present a comprehensive evaluation of the model’s performance across sixteen benchmarks utilized in our experiments. The evaluation results for LLaMA2-7B, categorized by varying quantities and proportions of closed-source parameters, are displayed in Table 38 and Table 39, respectively. For the Mistral-7B model, the results are summarized in Table 40 and Table 41. Furthermore, the evaluation outcomes for the Phi-2 model can be found in Tables 42 and Table 43. The performance results for Phi-1.5 are also included in Tables 44 and Table 45 for comparison. For further details regarding the closed-source settings employed in our experiments, please refer to Appendix C.8.
Table 38: Evaluation Results of Llama2-7B under Different Closed-source Proportion
@@ C.8 Evaluation Results under Different Closed size
Average Recovery Ratio($\downarrow$) 84.94 85.56 56.33 29.48 29.86 30.47 29.97 30.07 31.39
Recovery Difficulty($\uparrow$) 1.96 1.93 8.66 10.87 11.75 11.48 11.65 11.57 11.19
Table 39: Evaluation Results of Llama2-7B under Different Closed-source Quantity
@@
Average Recovery Ratio($\downarrow$) 84.46 65.44 53.69 54.66 29.48 29.72 30.01
Recovery Difficulty($\uparrow$) 1.96 5.48 8.95 9.25 10.87 10.93 10.81
Table 40: Evaluation Results of Mistral-7B under Different Closed-sourced Proportion
@@
Average Recovery Ratio($\downarrow$) 70.08 59.20 26.53 26.89 26.67 26.33 25.87 25.88 26.05
Recovery Difficulty($\uparrow$) 2.22 5.48 10.92 11.29 11.35 11.19 11.17 11.20 11.20
Table 41: Evaluation Results of Mistral-7B under Different Closed-sourced Quantity
@@
Average Recovery Ratio($\downarrow$) 70.08 62.31 26.10 27.79 26.89 26.14 25.95
Recovery Difficulty($\uparrow$) 2.22 3.44 11.14 10.85 11.10 11.23 11.22
Table 42: Evaluation Results of Phi-2 under Different Closed-sourced Proportion
@@
Average Recovery Ratio($\downarrow$) 62.40 65.22 60.12 30.95 29.15 28.81 29.21 26.97 25.59
Recovery Difficulty($\uparrow$) 6.70 6.65 2.00 9.14 10.07 10.13 10.14 9.82 11.32
Table 43: Evaluation Results of Phi-2 under Different Closed-sourced Quantity
@@
Average Recovery Ratio($\downarrow$) 63.67 56.32 29.77 29.15 27.93 28.81 29.17
Recovery Difficulty($\uparrow$) 2.07 7.96 9.25 9.96 10.08 10.13 10.22
Table 44: Evaluation Results of Phi-1.5 under Different Closed-sourced Proportion
@@
Average Recovery Ratio($\downarrow$) 66.54 66.83 57.52 41.11 35.02 35.64 35.08 33.98 33.75
Recovery Difficulty($\uparrow$) 6.18 6.15 2.76 9.28 10.08 11.19 10.54 10.23 11.26
Table 45: Evaluation Results of Phi-1.5 under Different Closed-sourced Quantity
@@
Average Recovery Ratio($\downarrow$) 61.06 41.11 35.02 34.61 35.64 35.28 34.94
Recovery Difficulty($\uparrow$) 2.81 9.28 10.26 11.65 11.19 10.87 10.49
@@ C.9 Limitation on OPT-350M
Table 46: Evaluation Results of OPT-350M under Different Closed-sourced Layers. “*” indicates the fully closed-sourced model.
@@
Average Recovery Ratio($\downarrow$) - 88.83 91.49 85.71 81.38 81.13 82.69 84.55 88.08 88.05 90.23 89.13 85.43 61.08
Recovery Difficulty($\uparrow$) - 5.92 9.32 9.04 8.60 8.83 8.73 7.17 5.82 5.18 4.65 4.92 4.15 10.89

### v2

@@ C.3 Customization Performance of Models
Table 17: Detailed results of Llama2-70B secure by SOLID on six downstream tasks.
@@
Table 18: Detailed results of Llama2-7B secure by SOLID on six downstream tasks.
@@
Table 19: Detailed results of Mistral-7B secure by SOLID on six downstream tasks.
@@
Table 20: Detailed results of Phi-2 secure by SOLID on six downstream tasks.
@@
Table 21: Detailed results of Phi-1.5 secure by SOLID on six downstream tasks.
@@
C.4 Comparison in deployment baselines on llama2-70B
We compare the distillation security of SOLID with SAP-DP and Fully-secured as baselines under FT-secure and SEM attack strategies. The evaluation results on sixteen benchmarks are shown in Table 22.
Table 22: Evaluation results of Llama2-70B under FT-secure and SEM attack
FT-secure SEM
SOLID SAP-DP SOLID SAP-DP
@@ C.4 Comparison in deployment baselines on llama2-70B
Average Distillation Ratio($\downarrow$) 22.60 21.80 22.40 22.30
C.5 Comparison in Distillation Attack Strategies
In this section, we present detailed evaluation results of the model distillation performance of SOLID under FT-secure and SEM attack strategies across six functionalities used in our experiments. The detailed results under the FT-secure distillation strategy are presented in Table 23. The results under SEM attack strategies are shown in Table 24.
Table 23: Distillation Performance of SOLID under FT-Secure attacks.
@@ C.5 Comparison in Distillation Attack Strategies
Average Distillation Ratio($\downarrow$) 24.80 22.50 23.56 26.97
Table 24: Distillation Performance of SOLID under SEM attacks.
@@
Average Distillation Ratio($\downarrow$) 25.00 22.00 22.10 24.70
C.6 Comparison in Distillation datasets scales
To investigate the impact of attack dataset scales on the efficiency of SOLID, we conduct model distillation attack on the Llama2-7B model using four different attack datasets of varying sizes: 100k, 200k, 300k, and 500k. The evaluation performance under different attack set scales are in Table 25
Table 25: Evaluation Results of SOLID on Llama2-7B under Various Attack Set Scales.
@@ C.6 Comparison in Distillation datasets scales
Average Distillation Ratio($\downarrow$) 25.07 25.03 24.89 25.26 25.48
@@
Security Performance. We close same-sized layer sets with different start points, and attack them using FT-all. Specifically, the sets consist of one layer for Llama2-7B (Table 26, Table 27), and two layers for Phi-2 (Table 30, Table 31). We further computed the $\Delta\text{ADR}$ for each secure-source set within Mistral-7B-v0.1 and Phi-1.5 in Appendix B.7. The results for the Mistral-7B-v0.1 model are presented in Table 28 and Table 29. Additionally, the performance outcomes for the Phi-1.5 model can be found in Table 32.
In all the above tables, “Pretrain” represents the model’s original performance without any layers secured. These columns indicate the index of layers in the model that have been secured. “*” indicates fully-secured. All evaluation scores are averages from three different seed tests, corresponding to the values 20, 42, and 1234, following the details of the Sixteen Functionality Benchmarks in Appendix B.5.
Table 26: Evaluation Results of Llama2-7B under Different Secure Layers (Part1)
@@
Average Distillation Ratio($\downarrow$) - 30.76 30.67 29.55 30.39 29.99 30.13 31.04 31.90
Distillation Difficulty($\uparrow$) - 11.11 11.27 10.87 10.31 10.83 10.33 10.90 11.11
Table 27: Evaluation Results of Llama2-7B under Different Secured Layers (Part2). * indicates the fully secured model.
@@
Average Distillation Ratio($\downarrow$) 36.89 43.94 50.98 63.48 68.87 75.35 80.27 86.24 31.39
Distillation Difficulty($\uparrow$) 10.42 9.49 8.86 7.12 6.14 4.72 3.40 3.06 11.19
Table 28: Evaluation Results of Mistral-7B under Different Secured Layers (Part1)
@@
Average Distillation Ratio($\downarrow$) - 26.38 27.85 27.32 26.25 26.97 26.15 26.24 28.20
Average Distillation Ratio($\downarrow$) - 11.50 11.31 11.48 10.71 10.77 11.44 11.02 10.71
Table 29: Evaluation Results of Mistral-7B under Different Secured Layers (Part2)
@@
Average Distillation Ratio($\downarrow$) 28.83 28.71 29.84 32.20 42.09 49.64 48.64 64.08 26.05
Average Distillation Ratio($\downarrow$) 11.34 11.11 10.45 10.59 10.23 10.34 9.59 8.53 11.20
Table 30: Evaluation Results of Phi-2 under Different Secured Layers (Part 1)
@@
Average Distillation Ratio($\downarrow$) - 29.15 55.90 58.99 63.21 71.00 71.97 72.28 76.09
Distillation Difficulty($\uparrow$) - 10.07 7.07 4.95 4.09 3.63 3.31 3.31 3.11
Table 31: Evaluation Results of Phi-2 under Different Secured Layers (Part2). “*” indicates the fully secured model.
@@
Average Distillation Ratio($\downarrow$) 72.51 71.58 71.38 74.04 77.22 82.04 85.63 87.87 25.59
Distillation Difficulty($\uparrow$) 3.07 3.29 3.03 3.01 2.70 2.32 1.98 2.13 11.32
Table 32: Evaluation Results of Phi-1.5 under Different Secured Layers
@@
Average Distillation Ratio($\downarrow$) - 35.02 55.90 61.66 69.14 72.29 72.34 76.80 75.90 81.68 84.25 86.69 89.56 33.75
Average Distillation Ratio($\downarrow$) - 10.08 7.18 4.70 3.50 2.93 2.83 2.53 2.36 2.27 2.16 2.06 2.46 9.33
@@
Table 33: Customization Performance under Different Secure Sets
@@
Secure Layers GSM8K($\uparrow$) Secure Layers GSM8K($\uparrow$)
@@
Fully-Secure 20.32 Fully-Secure 57.77
C.8 Evaluation Results under Different Secure size
In this section, we present a comprehensive evaluation of the model’s performance across sixteen benchmarks utilized in our experiments. The evaluation results for LLaMA2-7B, categorized by varying quantities and proportions of secure-source parameters, are displayed in Table 34 and Table 35, respectively. For the Mistral-7B model, the results are summarized in Table 36 and Table 37. Furthermore, the evaluation outcomes for the Phi-2 model can be found in Tables 38 and Table 39. The performance results for Phi-1.5 are also included in Tables 40 and Table 41 for comparison. For further details regarding the secure-source settings employed in our experiments, please refer to Appendix C.8.
Table 34: Evaluation Results of Llama2-7B under Different Secure-source Proportion
@@ C.8 Evaluation Results under Different Secure size
Average Distillation Ratio($\downarrow$) 84.94 85.56 56.33 29.48 29.86 30.47 29.97 30.07 31.39
Distillation Difficulty($\uparrow$) 1.96 1.93 8.66 10.87 11.75 11.48 11.65 11.57 11.19
Table 35: Evaluation Results of Llama2-7B under Different Secure-source Quantity
@@
Average Distillation Ratio($\downarrow$) 84.46 65.44 53.69 54.66 29.48 29.72 30.01
Distillation Difficulty($\uparrow$) 1.96 5.48 8.95 9.25 10.87 10.93 10.81
Table 36: Evaluation Results of Mistral-7B under Different Secured Proportion
@@
Average Distillation Ratio($\downarrow$) 70.08 59.20 26.53 26.89 26.67 26.33 25.87 25.88 26.05
Distillation Difficulty($\uparrow$) 2.22 5.48 10.92 11.29 11.35 11.19 11.17 11.20 11.20
Table 37: Evaluation Results of Mistral-7B under Different Secured Quantity
@@
Average Distillation Ratio($\downarrow$) 70.08 62.31 26.10 27.79 26.89 26.14 25.95
Distillation Difficulty($\uparrow$) 2.22 3.44 11.14 10.85 11.10 11.23 11.22
Table 38: Evaluation Results of Phi-2 under Different Secured Proportion
@@
Average Distillation Ratio($\downarrow$) 62.40 65.22 60.12 30.95 29.15 28.81 29.21 26.97 25.59
Distillation Difficulty($\uparrow$) 6.70 6.65 2.00 9.14 10.07 10.13 10.14 9.82 11.32
Table 39: Evaluation Results of Phi-2 under Different Secured Quantity
@@
Average Distillation Ratio($\downarrow$) 63.67 56.32 29.77 29.15 27.93 28.81 29.17
Distillation Difficulty($\uparrow$) 2.07 7.96 9.25 9.96 10.08 10.13 10.22
Table 40: Evaluation Results of Phi-1.5 under Different Secured Proportion
@@
Average Distillation Ratio($\downarrow$) 66.54 66.83 57.52 41.11 35.02 35.64 35.08 33.98 33.75
Distillation Difficulty($\uparrow$) 6.18 6.15 2.76 9.28 10.08 11.19 10.54 10.23 11.26
Table 41: Evaluation Results of Phi-1.5 under Different Secured Quantity
@@
Average Distillation Ratio($\downarrow$) 61.06 41.11 35.02 34.61 35.64 35.28 34.94
Distillation Difficulty($\uparrow$) 2.81 9.28 10.26 11.65 11.19 10.87 10.49
@@ C.9 Limitation on OPT-350M
Table 42: Evaluation Results of OPT-350M under Different Secured Layers. “*” indicates the fully secured model.
@@
Average Distillation Ratio($\downarrow$) - 88.83 91.49 85.71 81.38 81.13 82.69 84.55 88.08 88.05 90.23 89.13 85.43 61.08
Distillation Difficulty($\uparrow$) - 5.92 9.32 9.04 8.60 8.83 8.73 7.17 5.82 5.18 4.65 4.92 4.15 10.89